OpenStudy (anonymous):

find three consecutive integers whos sum is 240.

jimthompson5910 (jim_thompson5910):

let x the first integer x+1 is the next consecutive integer and x+2 is the next consecutive integer after the second integer

jimthompson5910 (jim_thompson5910):

the three consecutive integers must add to 240, so x+x+1+x+2 = 240

OpenStudy (anonymous):

240=x+(x+1)+(x+2)

OpenStudy (anonymous):

which simplifies to 240=3x+3 237=3x 79=x

OpenStudy (anonymous):

the integers are: 79, 80, and 81

OpenStudy (anonymous):

Hope this helped